# InversionView: A General-Purpose Method for Reading Information from Neural Activations

### v3

@@ Abstract
The inner workings of neural networks can be better understood if we can fully decipher the information encoded in neural activations. In this paper, we argue that this information is embodied by the subset of inputs that give rise to similar activations. Computing such subsets is nontrivial as the input space is exponentially large. We propose InversionView, which allows us to practically inspect this subset by sampling from a trained decoder model conditioned on activations. This helps uncover the information content of activation vectors, and facilitates understanding of the algorithms implemented by transformer models. We present four case studies where we investigate models ranging from small transformers to GPT-2. In these studies, we demonstrate the characteristics of our method, show the distinctive advantages it offers, and provide causally verified circuits. [^1] Code is available at https://github.com/huangxt39/InversionView
@@ 1 Introduction
Here, we introduce InversionView as a principled general-purpose method for generating hypotheses about the information present in activations in neural models on language and discrete sequences, which in turn helps us put together the algorithm implemented by the model. InversionView aims at providing a direct way of reading out the information encoded in an activation. The technique starts from the intuition that the information encoded in an activation can be formalized as its preimage, the set of inputs giving rise to this particular activation under the given model. In order to explore this preimage, given an activation, we train a decoder to sample from this preimage. Inspection of the preimage, across different inputs, makes it easy to identify which information is passed along, and which information is forgotten. It accounts for the geometry of the representation, and can identify which information is reinforced or downweighted at different model components. InversionView facilitates the interpretation workflow, and provides output that is in principle amenable to automated interpretation via LLMs (we present a proof of concept in Section 4).
We showcase the usefulness of the method in three thorough case studies: a character counting task, Indirect Object Identification, and 3-digit addition. We also present preliminary results on the factual recall task, demonstrating the applicability of our method to larger models. The character counting task illustrates how the method uncovers how information is processed and forgotten in a small transformer. In Indirect Object Identification in GPT2-Small [50], we use InversionView to easily interpret the information encoded in the components identified by [50], substantially simplifying the interpretability workflow. For 3-digit addition, we use InversionView to provide for the first time a fully verified circuit sufficient for performing the task. Across the case studies, InversionView allows us to rapidly generate hypotheses about the information encoded in each activation site. Coupled with attention patterns or patching methods, we reverse-engineer the flow of information, which we verify using causal interventions.
Figure 1: Illustration of the geometry at two different activation sites, encoding different information about the input. On the left, the semantics of being on leave are encoded. On the right, the information that the subject of the input sentence is John is encoded.
@@ 2.1 Interpretation Framework
2.2 Conditional Decoder Model
@@
We implement the decoder as an autoregressive language model conditioned on $\mathbf{z}^{q}$, decoding input samples ${\bf x}$. As the decoder’s training objective corresponds to recovering ${\bf x}$ exactly, sampling at temperature 1 will typically not cover the full $\epsilon$-preimage. Thus, for generating elements of the $\epsilon$-preimage, we increase diversity by drawing samples at higher temperatures and with noise added to $\mathbf{z}^{q}$ (details in Appendix A.2). Note that the activation $f({\bf x})$ will depend on the position in the input sequence; hence, we separately evaluate $D(f({\bf x}),{\bf z}^{q})$ at each position $i$ in each sample ${\bf x}$, and select the position $i$ minimizing $D(f({\bf x}),{\bf z}^{q})$,333 In some cases, including Figure 2(b) 3(b), we fix a particular position for interpretation. See Appendix A.3. determine membership in $B_{\mathbf{z}^{q},f,\epsilon}$, and subsample in-$\epsilon$-preimage and out-of-$\epsilon$-preimage samples for inspection.
An important question is whether this method, relying on a black-box decoder, produces valid $\epsilon$-preimages. We can approach this question from two perspectives, namely correctness (are all generated samples in the $\epsilon$-preimage?) and completeness (are the samples representative of the $\epsilon$-preimage?). As we evaluate $D(f({\bf x}),{\bf z}^{q})$ for each generated samples, correctness is ensured by design, thus we avoid making conclusion based on wrong samples. As for completeness, we ask whether all elements of the $\epsilon$-preimage can be output by the method at reasonable probability. If some groups of inputs in $\epsilon$-preimage are systematically missing from the generations, one may overestimate the information contained in activations. But this behavior would be punished by the training objective, since the loss on these examples would be high. We explicitly verify completeness by enumerating inputs in one of our case studies (Appendix B). Another approach is to design counter-examples ${\bf x}$ not satisfying a hypothesis about the content of $B_{\mathbf{z}^{q},f,\epsilon}$, and checking if they in fact are outside of it. In our experiments, we found that these examples were always outside of $B_{\mathbf{z}^{q},f,\epsilon}$.
@@ 3.1 Setup and Background
We train a single two-layer transformer decoder across all activation sites of interest; we chose this architecture for simplicity. The query activation ${\bf z}^{q}$ is concatenated with an activation site embedding ${\bf e}$, a learned embedding layer indicating where the activation comes from, passed through multiple MLP layers with residual connections, and then made available to the attention heads in each layer of the decoder, alongside the already present tokens from the input, so that each attention head can also attend to the post-processed query activation in addition to the context tokens. Each training example is a triple consisting of an activation vector ${\bf z}^{q}\in\mathbb{R}^{d}$, the activation site index, and the input, on which the decoder is trained with a language modeling objective. Appendix C has technical details.
@@ 3.2 Character Counting
We train a transformer (2 layers, 1 head) on inputs such as “vvzccvczvvvzvcvc||||v:8” to predict the last token “8”, the frequency of the target character (here, “v”) before the separator “||||”. For each input, three distinct characters are sampled from the set of lowercase characters, and each character’s frequency is sampled uniformly from 1–9. The input length varies between 7 and 31. We created 1.56M instances and applied a 75%-25% train-test split; test set accuracy is 99.53%. See details in Appendix D. For InversionView, we use $D(\mathbf{z},\mathbf{z}^{q})=\frac{\|\mathbf{z}-\mathbf{z}^{q}\|_{2}}{\|\mathbf{z}^{q}\|_{2}}$ (i.e., normalized euclidean distance), as the magnitude of activations varies between layers, and the threshold $\epsilon=0.1$.
Figure 2: InversionView on Character Counting Task. The model counts how often the target character (after ’|’) occurs in the prefix (before ’|’). B and E denote the beginning and end of sequence tokens. The query activation conditions the decoder to generate samples that capture its information content. We show non-cherrypicked samples inside and outside the $\epsilon$-preimage ($\epsilon=0.1$) at three activation sites on the same query input. Distance for each sample is calculated between activations corresponding to the parenthesized characters in the query input and the sample. True count indicates the correct count of the target character in the samples (decoder may generate an incorrect count). (a) MLP layer amplifies count information. By comparing the distances before (left) and after (right) the MLP, we can see samples with diverging counts become much more distant from the query activation. (b) In the next layer (when the information is copied from residual stream of target character to the residual stream of “:” via attention, since “:” exclusively attends to target character), the count is retained but the identity of the target character is no longer encoded (“c”, “m”, etc. instead of “g”), as it is no longer relevant for the predicting the count. Therefore, observing the generations corresponding to activation sites informs us of their content and how it changes across activation sites.
@@
In layer 0, the target character consistently attends to the same character in the previous context, suggesting that counting happens here. In Figure 2(a), we show the $\epsilon$-preimage of $x^{0,\mathrm{mid}}_{tc}$ and $x^{0,\mathrm{post}}_{tc}$, where the subscript ${tc}$ denotes the target character. We show $\approx 10$ random samples at a single query input, but our hypotheses are based on—and easily confirmed by—rapid visual inspection of dozens of inputs across different query inputs. [^4] Besides representative examples given in the Appendix, the reader can check generations conveniently at https://inversion-view.streamlit.app. On the left (before the MLP), the activation encodes the target character, as all samples have “g” as the target character. Count information is not sharply encoded: while the closest activation corresponds to “g” occurring 3 times, two activations corresponding to a count-4 input (“g” occurring 4 times) are also close, even closer than a count-3 input. On the other hand, on the right (after the MLP), only count-3 inputs are inside the $\epsilon$-preimage, and count-4 inputs become much more distant than before. Comparing the $\epsilon$-preimage before and after the MLP in layer 0, we find that the MLP makes the count information more prominent in the representational geometry of the activation. The examples are not cherry-picked; count information is generally reinforced by the MLP across query inputs.
In the next layer, the colon consistently attends to the target character, and InversionView confirms that count information is moved to the colon’s residual stream (Figure 2(b)). More importantly, this illustrates how information is abstracted: We previously found that $x^{0,\mathrm{post}}_{tc}$ encodes identity and frequency of the target character. However, the colon obtains only an abstracted version of the information, in which count information remains while the target character is largely (though not completely) removed. InversionView makes this process visible, by showing that the target character becomes interchangeable with little change to the activation. See more examples in Appendix D.2. Overall, with InversionView, we have found a simple algorithm by which the model makes the right prediction: In layer 0, the target character attends to all its occurrences and obtains the counts. In layer 1, the colon moves the results from the target character to its residual stream and then produces the correct prediction. Accounting for other activation sites, we find that the model implements a somewhat more nuanced algorithm, investigated in Appendix D.4. Overall, InversionView shows how certain information is amplified, providing an understanding of the effect of the MLP layer. More importantly, InversionView also shows how information is abstracted or forgotten.
@@ 3.3 IOI circuit in GPT-2 small
We start with the Name Mover Head 9.9, which Wang et al. [50] found moves the IO name to the residual stream of END. 3(b) shows the $\epsilon$-preimage at “to”. The samples in the $\epsilon$-preimage share the name “Justin” as the IO. The head also shows similar activity at some other positions (Appendix A.3). Results are consistent across query inputs. Therefore, the result of InversionView is consistent with the conclusions of Wang et al. [50] on head 9.9. Applying the same analysis to other heads (Table 2), we recovered information in high agreement with the information that [50] had inferred using multiple tailored methods. For example, Wang et al. [50] found S-Inhibition heads were outputting both token signals (value of S) and position signals (position of S1) by patching these heads’ outputs from a series of counterfactual datasets. These datasets are designed to disentangle the two effects, in which token and/or position information are ablated or inverted. On the other hand, these two kinds of information can be directly read out by InversionView (Figure 18 shows an example for one of the S-Inhibition heads that contains position information), and there is no need to guess the possible information to design patching experiments. Overall, among the 26 attention heads that Wang et al. [50] identified, InversionView indicates a different interpretation in only 3 cases; these (0.1, 0.10, 5.9) were challenging for the methods used before (Appendix E.3). In summary, InversionView scales to larger models.
@@
Figure 5(a) shows the circuit predicting A1. InversionView allows us to diagnose an important deficiency of this circuit: even though the ones place sometimes receives some attention in layer 1, the circuit does not consistently provide the carry from the ones place to the hundreds place, which matters on certain inputs—we find that this deficiency in the circuit accounts for all mistakes made by the model (Appendix F.3). Taken together, we have provided a circuit allowing the model to predict A1 while also understanding its occasional failure in doing so correctly. Corresponding findings for A2, A3, and A4 are in Table 3 and Figure 30. From A2 onwards, InversionView allows us to uncover how the model exhibits two different algorithms depending on whether the resulting output will have 3 or 4 digits. In particular, when predicting A3, the layer 0 circuit is the same across both cases, while the layer 1 circuit varies, since this determines whether A3 will be a tens place or ones place. Beyond figures in the Appendix, we also encourage readers to verify our claims in our interactive web application.
@@
Figure 6: 3-Digit Addition Task: InversionView uncovers different ways in which digit is encoded in activations. Left: The digits in the hundreds place are encoded separately and hence generations denote them as separate entities. Right: The digits in the tens place are encoded as a sum (9 in this case) and the generations represent different 2-partitions (7+2, 6+3, 1+8, 5+4, etc.) of that sum.
@@
Heads often read from both digits of the same place. Only the sum matters for addition – raising the question whether the digits are represented separately, or only as their sum. Unlike traditional probing, InversionView answers this question without designing any tailored probing tasks. In Figure 6 (left), $a^{0,2}$ exactly represents F2 and S2 (here, 2 and 5). Other inputs where F1+S1=5+2 have high $D$. In contrast, on the right, F2 and S2 are represented only by their sum: throughout the $\epsilon$-preimage, F2+S2=9. In fact, we find such sum-only encoding only when F2+S2=9—a special case where the ones place of operands affects the hundreds place of the answer via cascading carry. We hypothesize that the model encodes them similarly because these inputs require special treatment. Therefore, even though encoding number pairs by their sum is a good strategy for the addition task from a human perspective, the model only does it as needed. We also observe intermediate cases (Figure 28).

### v4

@@ Abstract
The inner workings of neural networks can be better understood if we can fully decipher the information encoded in neural activations. In this paper, we argue that this information is embodied by the subset of inputs that give rise to similar activations. We propose InversionView, which allows us to practically inspect this subset by sampling from a trained decoder model conditioned on activations. This helps uncover the information content of activation vectors, and facilitates understanding of the algorithms implemented by transformer models. We present four case studies where we investigate models ranging from small transformers to GPT-2. In these studies, we show that InversionView can reveal clear information contained in activations, including basic information about tokens appearing in the context, as well as more complex information, such as the count of certain tokens, their relative positions, and abstract knowledge about the subject. We also provide causally verified circuits to confirm the decoded information. [^1] Code is available at https://github.com/huangxt39/InversionView
@@ 1 Introduction
Here, we introduce InversionView as a principled general-purpose method for generating hypotheses about the information present in activations in neural models on language and discrete sequences, which in turn helps us identify how the information flows through the model—crucial for obtaining the algorithm implemented by the model. InversionView aims at providing a direct way of reading out the information encoded in an activation. The technique starts from the intuition that the information encoded in an activation can be formalized as its preimage, the set of inputs giving rise to this particular activation under the given model. In order to explore this preimage, given an activation, we train a decoder to sample from this preimage. Inspection of the preimage, across different inputs, makes it easy to identify which information is passed along, and which information is forgotten. It accounts for the geometry of the representation, and can identify which information is reinforced or downweighted at different model components. InversionView facilitates the interpretation workflow, and provides output that is in principle amenable to automated interpretation via LLMs (we present a proof of concept in Section 4).
We showcase the usefulness of the method in three case studies: a character counting task, Indirect Object Identification, and 3-digit addition. We also present preliminary results on the factual recall task, demonstrating the applicability of our method to larger models. The character counting task illustrates how the method uncovers how information is processed and forgotten in a small transformer. In Indirect Object Identification in GPT2-Small [54], we use InversionView to easily interpret the information encoded in the components identified by Wang et al. [54], substantially simplifying the interpretability workflow. For 3-digit addition, we use InversionView to provide for the first time a fully verified circuit. Across the case studies, InversionView allows us to rapidly generate hypotheses about the information encoded in each activation site. Coupled with attention patterns or patching methods, we reverse-engineer the flow of information, which we verify using causal interventions.
Figure 1: Illustration of the geometry at two different activation sites, encoding different information about the input. Top: the semantics of being on leave are encoded. Bottom: the information that the subject of the input sentence is John is encoded.
@@ 2 Methodology
Figure 2: (a) The probed model is trained on language modeling objective. (b) Given a trained probed model, we first cache the internal activations $\mathbf{z}$ together with their corresponding inputs and activation site indices (omitted in the figure for brevity), then use them to train the decoder. The decoder is trained with language modeling objective, while being able to attend to $\mathbf{z}$. (c) When interpreting a specific query activation $\mathbf{z^{q}}$, we give it to the decoder, which generates possible inputs auto-regressively. We then evaluate the distances on the original probed model.
Conditional Decoder Model
@@
We implement the decoder as an autoregressive language model conditioned on $\mathbf{z}^{q}$, decoding input samples ${\bf x}$ (see Figure 2, and details in Appendix C). As the decoder’s training objective corresponds to recovering ${\bf x}$ exactly, sampling at temperature 1 will typically not cover the full $\epsilon$-preimage. Thus, for generating elements of the $\epsilon$-preimage, we increase diversity by drawing samples at higher temperatures and with noise added to $\mathbf{z}^{q}$ (details in Appendix A.2). Note that the activation $f({\bf x})$ will depend on the position in the input sequence; hence, we separately evaluate $D(f({\bf x}),{\bf z}^{q})$ at each position $i$ in each sample ${\bf x}$, and select the position $i$ minimizing $D(f({\bf x}),{\bf z}^{q})$,333 In some cases, including Figure 3(b) 4(b), we fix a particular position for interpretation. See Appendix A.3. determine membership in $B_{\mathbf{z}^{q},f,\epsilon}$, and subsample in-$\epsilon$-preimage and out-of-$\epsilon$-preimage samples for inspection.
An important question is whether this method, relying on a black-box decoder, produces valid $\epsilon$-preimages. Correctness (are all generated samples in the $\epsilon$-preimage?) is ensured by design, as we evaluate $D(f({\bf x}),{\bf z}^{q})$ for each generated sample. The other angle is completeness (are the samples representative of the $\epsilon$-preimage?). If some groups of inputs in $\epsilon$-preimage are systematically missing from the generations, one may overestimate the information contained in activations. But this behavior would be punished by the training objective, since the loss on these examples would be high. We explicitly verify completeness by enumerating inputs in one of our case studies (Appendix B). Another approach is to design counter-examples ${\bf x}$ not satisfying a hypothesis about the content of $B_{\mathbf{z}^{q},f,\epsilon}$. In our experiments, we found that these examples were always outside of $B_{\mathbf{z}^{q},f,\epsilon}$.
Figure 3: InversionView on Character Counting Task. The model counts how often the target character (after ’|’) occurs in the prefix (before ’|’). B and E denote beginning and end of sequence tokens. The query activation conditions the decoder to generate samples capturing its information content. We show non-cherrypicked samples inside and outside the $\epsilon$-preimage ($\epsilon=0.1$) at three activation sites on the same query input. Distance for each sample is calculated between activations corresponding to the parenthesized characters in the query input and the sample. “True count” indicates the correct count of the target character in the samples (decoder may generate incorrect counts). (a) MLP layer amplifies count information. Comparing the distances before (left) and after (right) the MLP, we see that samples with diverging counts become much more distant from the query activation. (b) In the next layer (“:” exclusively attends to target character – copying information from residual stream of target character to the residual stream of “:”), the count is retained but the identity of the target character is no longer encoded (“c”, “m”, etc. instead of “g”), as it is no longer relevant for the predicting the count. Therefore, observing the generations informs us of the activations’ content and how it changes across activation sites.
@@ 3 Discovering the Underlying Algorithm by InversionView
We train a single two-layer transformer decoder across all activation sites of interest. The query activation ${\bf z}^{q}$ is concatenated with an activation site embedding ${\bf e}$, a learned embedding layer indicating where the activation comes from, passed through multiple MLP layers with residual connections, and then made available to the attention heads in each layer of the decoder, alongside the already present tokens from the input, so that each attention head can also attend to the post-processed query activation in addition to the context tokens. Each training example is a triple consisting of an activation vector ${\bf z}^{q}\in\mathbb{R}^{d}$, the activation site index, and the input, on which the decoder is trained with a language modeling objective. Appendix C has technical details.
@@ 3.1 Character Counting
We train a transformer (2 layers, 1 head) on inputs such as “vvzccvczvvvzvcvc||||v:8” to predict the last token “8”, the frequency of the target character (here, “v”) before the separator “||||”. For each input, three distinct characters are sampled from the set of lowercase characters, and each character’s frequency is sampled uniformly from 1–9. The input length varies between 7 and 31. We created 1.56M instances and applied a 75%-25% train-test split; test set accuracy is 99.53% (Details in Appendix D). We use $D(\mathbf{z},\mathbf{z}^{q})=\frac{\|\mathbf{z}-\mathbf{z}^{q}\|_{2}}{\|\mathbf{z}^{q}\|_{2}}$ (i.e., normalized euclidean distance, as the magnitude of activations varies between layers), where $\mathbf{z}$ denotes the aforementioned $f(\mathbf{x})$, and set $\epsilon=0.1$.
@@
In layer 0, the target character consistently attends to the same character in the previous context, suggesting that counting happens here. In Figure 3(a), we show the $\epsilon$-preimage of $x^{0,\mathrm{mid}}_{tc}$ and $x^{0,\mathrm{post}}_{tc}$, where the subscript ${tc}$ denotes the target character. We show $\approx 10$ random samples at a single query input, but our hypotheses are based on—and easily confirmed by—rapid visual inspection of dozens of inputs across different query inputs.444 The reader can check generations conveniently at https://inversion-view.streamlit.app. On the left (before the MLP), the activation encodes the target character, as all samples have “g” as the target character. Count information is not sharply encoded: while the closest activation corresponds to “g” occurring 3 times, two activations corresponding to a count-4 input (“g” occurring 4 times) are also close, even closer than a count-3 input. On the other hand, on the right (after the MLP), only count-3 inputs are inside the $\epsilon$-preimage, and count-4 inputs become much more distant than before. Comparing the $\epsilon$-preimage before and after the MLP in layer 0, we find that the MLP makes the count information more prominent in the representational geometry of the activation. The examples are not cherry-picked; count information is generally reinforced by the MLP across query inputs.
In the next layer, the colon consistently attends to the target character, and InversionView confirms that count information is moved to the colon’s residual stream (Figure 3(b)). More importantly, this illustrates how information is abstracted: We previously found that $x^{0,\mathrm{post}}_{tc}$ encodes identity and frequency of the target character. However, the colon obtains only an abstracted version of the information, in which count information remains while the target character is largely (though not completely) removed. InversionView makes this process visible, by showing that the target character becomes interchangeable with little change to the activation. See more examples in Appendix D.2. Overall, with InversionView, we have found a simple algorithm by which the model makes the right prediction: In layer 0, the target character attends to all its occurrences and obtains the counts. In layer 1, the colon moves the results from the target character to its residual stream and then produces the correct prediction. Accounting for other activation sites, we find that the model implements a somewhat more nuanced algorithm, investigated in Appendix D.4. Overall, InversionView shows how certain information is amplified, but also how information is abstracted or forgotten.
@@ 3.2 IOI circuit in GPT-2 small
We start with the Name Mover Head 9.9, which Wang et al. [54] found moves the IO name to the residual stream of END. 4(b) shows the $\epsilon$-preimage at “to”. The samples in the $\epsilon$-preimage share the name “Justin” as the IO. The head also shows similar activity at some other positions (Appendix A.3). Results are consistent across query inputs. Therefore, InversionView agrees with the conclusions of Wang et al. [54] on head 9.9. Applying the same analysis to other heads (Table 2), we recovered information in high agreement with the information that Wang et al. [54] had inferred using multiple tailored methods. For example, Wang et al. [54] found S-Inhibition heads were outputting both token signals (value of S) and position signals (position of S1) by patching these heads’ outputs from a series of counterfactual datasets. These datasets are designed to disentangle the two effects, in which token and/or position information are ablated or inverted. These two kinds of information can be directly read out by InversionView (Figure 19 shows an example for an S-Inhibition head that contains position information), and there is no need to guess the possible information to design patching experiments. Overall, among the 26 attention heads that Wang et al. [54] identified, InversionView indicates a different interpretation in only 3 cases; these (0.1, 0.10, 5.9) were challenging for the methods used before (Appendix E.3). In summary, InversionView scales to larger models.
@@
Figure 6(a) shows the circuit predicting A1. InversionView allows us to diagnose an important deficiency of this circuit: Even though the ones place sometimes receives attention in layer 1, the circuit does not consistently provide the carry from the ones place to the hundreds place, which matters on certain inputs—we find that this deficiency in the circuit accounts for all mistakes made by the model (Appendix F.3). Taken together, we have provided a circuit allowing the model to predict A1 while also understanding its occasional failure in doing so correctly. Corresponding findings for A2, A3, and A4 are in Table 3 and Figure 31. From A2 onwards, InversionView allows us to uncover how the model exhibits two different algorithms depending on whether the resulting output will have 3 or 4 digits. In particular, when predicting A3, the layer 0 circuit is the same across both cases, while the layer 1 circuit varies, since this determines whether A3 will be a tens place or ones place. Beyond figures in the Appendix, we also encourage readers to verify our claims in our interactive web application.
@@
Figure 7: 3-Digit Addition Task: InversionView uncovers different ways in which digit representation is encoded in activations. Left: The digits in the hundreds place are encoded separately and hence generations denote them as separate entities. Right: The digits in the tens place are encoded as a sum (9 in this case) and the generations represent different 2-partitions (7+2, 6+3, 1+8, 5+4, etc.) of that sum.
@@
Heads often read from both digits of a place, but only the sum matters for addition. Are the digits represented separately, or only as their sum? Unlike traditional probing, InversionView answers this question without designing tailored probing tasks. In Figure 7 (left), $a^{0,2}$ exactly represents F2 and S2 (here, 2 and 5). Other inputs where F1+S1=5+2 have high $D$. In contrast, on the right, F2 and S2 are represented only by their sum: throughout the $\epsilon$-preimage, F2+S2=9. In fact, we find such sum-only encoding only when F2+S2=9—a special case where the ones place of operands affects the hundreds place of the answer via cascading carry. We hypothesize that the model encodes them similarly because these inputs require special treatment. Therefore, even though encoding number pairs by their sum is a good strategy for the addition task from a human perspective, the model only does it as needed. We also observe intermediate cases (Figure 29).